(1 point) (1 point) Find dine and at the given point without eliminating the parameter. 1 = {+2, y= +* +5t, t= 2.
Answer #1

E Queffrin * j*+2 ; 95 +54 ; *=2 > x=42+2 differensial it with respect to ® : 11 2 * . 39= **+5% differensiat it with respectNaw 5+ ك | ل ) ا dk الوزراء * ... - سلام . . . . 2 - 5 3 at t = 2 335 (=( ه )